### v1

@@ 4 Summary and Conclusions
Of course, our observations are specific to the 2HDM that we have considered here. Nonetheless, we can expect our findings to generalise to other scenarios as well, qualitatively: Moving away from the SM to address (parts of) its shortcomings in the light of the current consistency of LHC results with SM expectations will leave sizeable quantum footprints that can be targeted at experimental precision environments. In this sense, our results are representative and have provided convincing evidence that such parameter regions can be efficiently probed by $e^{+}e^{-}\to hZ$ measurements, which are driven by relatively model-independent $Z$-recoil studies. Therefore, if the more profound theory that underpins the SM follows the behaviour of established quantum field theory, there is no a priori reason beyond tuning for it to be missed at the next generation of lepton colliders at the latest.
Finally, it is worthwhile mentioning that the new physics as considered here can be discovered at any stage in the present and future collider programme, e.g. through small coupling deviations of the SM-like Higgs boson at the HL-LHC in the near future (e.g. through improving beyond the single-Higgs sensitivity expectations with growing data sets Belvedere:2024wzg or through discovery in rare channels). In parallel, with experiments actively including interference effects in their exotics’ searches for $gg\to H/A$, a coupling deviation at the HL-LHC could be supported by evidence for exotics at the LHC. This also includes search channels that we have not considered in this work. In such an instance, the combination of a $Z$ pole programme in tandem with an experimentally robust precision investigation of Higgs production at an electron-positron collider will lead to a detailed understanding of the nature of the electroweak scale.

### v3

@@ 4 Summary and Conclusions
Of course, our observations are specific to the 2HDM that we have considered here. Nonetheless, we can expect our findings to generalise to other scenarios as well, qualitatively. Moving away from the SM to address (parts of) its shortcomings in the light of the current consistency of LHC results with SM expectations will leave sizeable quantum footprints that can be targeted at experimental precision environments. In this sense, our results are representative and have provided convincing evidence that such parameter regions can be efficiently probed by $e^{+}e^{-}\to hZ$ measurements, which are driven by relatively model-independent $Z$-recoil studies. Therefore, if the more profound theory that underpins the SM follows the behaviour of established quantum field theory, there is no a priori reason beyond tuning for it to be missed at the next generation of lepton colliders at the latest.
Along this line, it is worthwhile pointing out that other $e^{+}e^{-}$ machines have been proposed, with or without a $Z$ pole programme, e.g. ILC Bambade:2019fyw , CLIC CLICdp:2018cto or LCF LinearColliderVision:2025hlt ; LinearCollider:2025lya . The precision of an $hZ$ measurement will be quantitatively similar. A less precise measurement of the electroweak input parameters that is provided by a dedicated $Z$ run can result in somewhat larger theoretical interpretation uncertainty. We have not considered this in detail, and within the approximations of our work, we can expect such a machine to have a similar sensitivity reach as the FCC-ee concept. Relaxing the $S,T,U$ constraints, which serve the purpose of aligning the Higgs couplings alongside the exotics’ mass splitting, a more general phenomenological NLO pattern becomes possible. The additional parametric freedom can then also lead to NLO cross sections consistent with the SM uncertainty. Such a limitation can potentially be overcome by studying exclusive decay modes in $Zh$ production, which provide a largely independent measurement of couplings and mixing angles as the Higgs boson is narrow.
Finally, it is worthwhile mentioning that the new physics as considered here can be discovered at any stage in the present and future collider programme, e.g. through small coupling deviations of the SM-like Higgs boson at the HL-LHC in the near future (e.g. through improving beyond the single-Higgs sensitivity expectations with growing data sets Belvedere:2024wzg or through discovery in rare channels). We highlight again that the experiments are actively including interference effects in their exotics’ searches for $gg\to H/A$, and we can expect analyses to fully reflect model correlations beyond benchmarks, especially when we turn to the HL-LHC phase, when backgrounds will become increasingly under statistical and systematic control. This also includes search channels that we have explicitly not considered in this work. In the lucky instance of new physics discovery, the combination of a $Z$ pole programme (for a more comprehensive analysis see Refs. Chen:2018shg ; Chen:2019pkq ) in tandem with an experimentally robust precision investigation of Higgs production at an electron-positron collider will lead to a detailed understanding of the nature of the electroweak scale. Whether or not an $e^{+}e^{-}$ collider remains a motivated concept will, of course, depend on the nature of the discovery, which could instead justify a direct consideration of FCC-hh.